The interior angles of polygon are in A.P. If the smallest angle is 120 degree and the common different is 5, then the number of sides of the polygon is :

Grade:11

1 Answers

Rounak banerjee
7 years ago
Sum of all int. angles = 180( n - 2)first term(a)=120d=5sum=n/2(2a+(n-1)d)so n(240+5n-5)=360(n-2) Solve for n